OpenStudy (anonymous):

In her recipe for homemade pasta sauce, Becky uses green peppers and tomato sauce in a ratio of 1 to 3. If she uses 6 cans of tomato sauce, how many green peppers should she use?

OpenStudy (anonymous):

helppppppppp pleaseeeeeeee

OpenStudy (anonymous):

i ll make it the best response

OpenStudy (ghazi):

1/3= cans of green pepper/ cans of sauce= cans/6\[\frac{ 1 }{ 3 }=\frac{ cans }{ 6 }\] solve for cans :)

OpenStudy (anonymous):

thanks~!

OpenStudy (ghazi):

:)